# The Need for Benchmarks to Advance AI-Enabled Player Risk Detection in Gambling

### v1

@@ 5.6 The Evolving Gambling Landscape and need for Dynamic Benchmarks
A final challenge is the dynamic and evolving nature of gambling. Models built to accurately classify or predict stable outcomes such as handwriting or images can be continuously evaluated against benchmark datasets with little need for updating. By contrast, gambling environments and consumer behavior can be continuously impacted by changes in regulation, local economies, and product innovation, which may result in data drift. Validation of a model against a historical dataset (such as those added to The Transparency Project) may hold little weight when evaluating contemporary performance. As such, ML benchmarks for gambling cannot be static. Government mandates that require data sharing (see Newall2024) are likely to be the most successful way of accessing periodically updated datasets for a benchmarking suite. Regardless, successfully sharing and deploying the first iteration of benchmarking datasets without compromising privacy will be essential to building trust with operators and the public alike, and ensuring continuous buy-in from industry stakeholders.
@@ 6 Conclusion
While we outline a broad range of challenges and considerations, we argue that these are not insurmountable and can be addressed with the right combination of data access, domain and technical expertise, and the ability to develop trustworthy solutions. A failure of industry to address these concerns will cast doubt on the efficacy of these solutions and increase pressure on regulators to take greater control of how they are developed and used. In Spain, the Directorate General for the Regulation of Gambling (DGOJ) recently received a legal mandate to establish a unified ML model that will be mandatory for all licensed operators in Spain (directorate2025). Such regulatory developments are likely to continue expanding if the transparency of existing industry methods remains opaque.
ORCID iDs
Disclosures
During the last five years, International Gaming Institute (IGI) at University of Nevada, Las Vegas, has received funding for its research and programs from Action Gaming, American Gaming Association, Aristocrat Leisure Limited, Association of Gaming Equipment Manufacturers, Axes.ai, Bet Blocker, Clarion Gaming, DraftKings, Entain Foundation, ESPN, Evoke plc, Focal Research Consultants, Gaming Analytics, Global Payments, IGT, Kindbridge Behavioral Health, Las Vegas Sands Corporation, Massachusetts Gaming Commission, MGM Resorts International, Playtech plc, Responsible Online Gaming Association, Yuhaaviatam of San Manuel Nation, Sightline Payments, Sports Betting Alliance, State of Nevada Department of Health and Human Services, State of Nevada Knowledge Fund, Walker Digital Table Systems, and Wynn Resorts Ltd. Additionally, IGI organizes the triennial International Conference on Gambling and Risk Taking, a research-focused event supported by sponsors from industry, academia, and the legal/regulatory sectors; a full list of sponsors for the most recent conference can be found at https://www.unlv.edu/igi/conference/18th/sponsors. IGI is home to an industry-focused advisory board (https://www.unlv.edu/igi/advisory-board), and specific programs, such as AiR Hub, have their own advisory panels. These advisory roles include resource support, and individual advisors are required to adhere to IGI policies. IGI maintains a strict research policy (https://www.unlv.edu/igi/research-policy), as well as a partnership and transparency framework (https://www.unlv.edu/igi/policies/partnership), to ensure appropriate firewalls exist between funding entities and IGI’s research and programs.
During the past 5 years, Kasra Ghaharian has received funding for research and/or consulting services from the Nevada Department of Health and Human Services, the Nevada Governor’s Office of Economic Development, the Massachusetts Gaming Commission, AXES.ai, Playtech, Sightline, IGT, Differential, Focal Research Consultants, GP Consulting, and the International Center for Responsible Gaming. Ghaharian has received honoraria/travel reimbursement from the Responsible Gambling Council, the Illinois Council on Problem Gambling, and Kindred Group. None of these entities played roles in the design, analysis, or interpretation of research, and imposed no constraints on publishing.
During the past 5 years Simo Dragicevic has received funding for employment, research and consulting services from The European Lotteries, The Gambling Commission (Great Britain), Playtech Plc, and Yaspa. Simo is also Adjunct Fellow at the International Gaming Institute (IGI).
Chris Percy provides consulting services to the gambling industry, with recent activity including research for the UK Gambling Commission and the BetBuddy player risk detection software and responsible gambling R&D at Playtech Plc.
During the past five years, Sarah E. Nelson has served as a paid grant reviewer for the New Frontiers in Research Fund (Government of Canada), the International Center for Responsible Gaming (ICRG), and the National Institutes of Health (NIH). She has also received travel reimbursement and speaker honoraria from the ICRG, Responsible Gaming Association of New Mexico, Responsibility.org, and the Maryland Highway Safety Office. She received publication royalty fees from the American Psychological Association, and received course royalty fees from the Harvard Medical School Department of Continuing Education.
W. Spencer Murch holds a postdoctoral fellowship from the Alberta Gambling Research Institute, a joint initiative by the Alberta Provincial Government, and the Universities of Calgary, Lethbridge, and Alberta. WSM holds a research grant from the International Centre for Responsible Gaming (USA) that did not fund his contributions to this research. WSM served on the Advisory Board on Safer Gambling, an expert group that formerly advised the UK Gambling Commission.
Robert M. Heirene’s contribution to this work was supported by a post-doctoral fellowship from the New South Wales Responsible Gambling Fund as part of the Gambling Research Capacity Grant program, administered by the Office of Responsible Gambling. RMH has worked on a project funded by Responsible Wagering Australia (a representative body of Australian online wagering operators; University of Sydney, 2019–2021) and as an independent, sub-contracted statistical consultant for PRET Solutions Inc on a commissioned project (funded by the Australian Casino operator Crown; 2023). RMH has also received research funding from the International Centre for Responsible Gaming (ICRG) and the Brain and Mind Center, University of Sydney.
Khalil Simeon-Rose declares the following financial interests/personal relationships which may be considered as potential competing interests: In the three-year period from 2022 to 2025, KS received research funding from the Cambridge Health Alliance, Division on Addiction; he received consulting payments from Eilers & Krejcik Gaming, Scientific Affairs, VictorStrategies, MGM Resorts, and the United Arab Emirates’ General Commercial Gaming Regulatory Authority; and he received expert witness payments from Benjamin Leigh Carter. KS holds unpaid appointments on the advisory board of the Responsible Online Gaming Association, the Canadian Gaming Association’s responsible gaming sub-committee, and the Deutsche Stiftung Glücksspielforschung gGmbH Scientific Advisory Board.
Tracy Schrans declares that, during the past 5 years, Focal Research Consultants have received funding from various gaming companies, regulatory bodies, government funding agencies, and research institutions for consultancy work in the area of player data analytics, social responsibility, and safer gambling products, services, and compliance solutions. None of these entities played a role in the design, analysis, or interpretation of this study, and imposed no constraints on publishing.

### v2

@@ 5.6 The Evolving Gambling Landscape and need for Dynamic Benchmarks
A final challenge is the dynamic and evolving nature of gambling. Models built to accurately classify or predict stable outcomes such as handwriting or images can be continuously evaluated against benchmark datasets with little need for updating. By contrast, gambling environments and consumer behavior can be continuously impacted by changes in regulation, local economies, and product innovation, which may result in data drift. Validation of a model against a historical dataset (such as those added to The Transparency Project) may hold little weight when evaluating contemporary performance. As such, ML benchmarks for gambling cannot be static. Government mandates that require data sharing (see Newall and Swanton, 2024) are likely to be the most successful way of accessing periodically updated datasets for a benchmarking suite. Regardless, successfully sharing and deploying the first iteration of benchmarking datasets without compromising privacy will be essential to building trust with operators and the public alike, and ensuring continuous buy-in from industry stakeholders.
@@ 6 Conclusion
While we outline a broad range of challenges and considerations, we argue that these are not insurmountable and can be addressed with the right combination of data access, domain and technical expertise, and the ability to develop trustworthy solutions. A failure of industry to address these concerns will cast doubt on the efficacy of these solutions and increase pressure on regulators to take greater control of how they are developed and used. In Spain, the Directorate General for the Regulation of Gambling (DGOJ) recently received a legal mandate to establish a unified ML model that will be mandatory for all licensed operators in Spain (Directorate General for the Regulation of Gambling, 2025). Such regulatory developments are likely to continue expanding if the transparency of existing industry methods remains opaque.